# Holes in Latent Space: Topological Signatures Under Adversarial Influence

### v1

@@ 4.1 Global Results
PCA and CCA. Figure 7 presents PCA results of the pruned barcode summaries, showing a clear separation between clean and poisoned subsamples across layers, which is consistent with the motivating experiment on a single subsample of clean vs. poisoned activations (Figure 1). We then tested the impact of each individual feature on the appearance of this separation. The projection of features onto the first principal component in Figure 7 shows that the mean of the deaths of 0-bars was the dominant contributor for layers 1, 8, 16, and 24. For layer 32, it was the second most significant contributor, with the number of 1-bars taking precedence as the primary contributor. The number of 1-bars also ranked as the second most significant feature for layers 24 and 16. In contrast, for layer 8, the second highest contribution came from the minimum of the deaths of 0-bars, while for layer 1, it was the standard deviation of the deaths of 0-bars. This observation is supported by a CCA between the pruned barcode summaries and the principal components of the PCA. CCA is a statistical method that quantifies linear relationships between two multivariate datasets by finding pairs of canonical variables with maximal correlation. The loadings are the contributions of individual features to these canonical variables, measuring their importance in capturing the relationship. We again found that mean of the deaths of the 0-bars ranked first in all layers, and that the number of 1-bars appeared as a significant statistic as well (see Figure 16).
@@
Figure 7: PCA of pruned barcode summaries of clean vs. poisoned activations. Clear distinction appears in the projection onto the two first principal components from the PCA of the pruned barcode summaries for layers 1, 8, 16, 24, and 32. The explained variance is 0.59, 0.49, 0.52, 0.96 and 0.83, respectively.
Regression and SHAP Analysis. We trained a logistic regression on the pruned barcode summaries, with a 70/30 split between train and test. The results of the regression plotted in the PCA projection, for visualization purposes, can be found in Figure 8. We obtained perfect accuracy and AUC–ROC, when testing on the test data, and 5-fold cross validation over the training data for all models. We used Shapley (or SHAP) values to interpret the exceptional performance of the regression model. Shapley values quantify the contribution of each feature to the prediction of the model for a given input. Figure 9 shows beeswarm plots of Shapley values for layers 1, 8, 16, 24 and 32, where each row represents a feature, and points correspond to the SHAP values of the input data (spread across the $x$-axis), colored by feature value for the corresponding input data point. Our analysis reveals that the mean of 0-bar deaths and the number of 1-bars strongly influence predictions, exhibiting a clear dichotomous effect.
@@
Figure 8: Logistic regression for clean vs. poisoned activations trained on a 70/30 train/test split of the pruned barcode summaries, plotted on the projection onto the two first PCs. Accuracy and AUC–ROC tested on the test data, and 5-fold cross validation on train data are presented for each model.
{adjustwidth}
-0.75in-0.25in
@@
Interpretation: The Shape of Adversarial Influence. Interpreting the distributions of the barcode summaries for clean vs. poisoned data reveals that adversarial conditions typically yield fewer 1-bars (loops) forming at later scales, yet persisting longer (see Figure 17). Conversely, the non-adversarial conditions tend to form earlier loops with more uniform lifetimes (higher persistent entropy). This pattern aligns with the Shapley value results (Figure 9): lower mean death times of 0-bars (i.e., more compact point clouds) are associated with predictions of “clean”, while higher values (more spread-out clouds) shift predictions toward “poisoned”. Similarly, a lower number of 1-bars tends to indicate “poisoned”, whereas a higher count suggests “clean”. In addition, a local dispersion ratio (Appendix A.2) and average cosine distance (Figure A.3) substantiated these results, revealing that adversarially influenced representation vectors become more dispersed (higher cosine distance) or concentrate variance onto specific axes (leading to flips in dispersion ratio), implying a reallocation of representational capacity toward a smaller number of large-scale features. In contrast, the non-adversarial conditions produced lower or more stable distance measures (less reconfiguration in the hidden space). Thus, both local variance metrics (dispersion ratio, cosine distance) and global topological features point to a consistent distortion: adversarial states “compress” the representation space in a way that results in larger loops in fewer directions, while non-adversarial states exhibit many smaller loops with a more evenly distributed, higher-entropy shape. See Appendix A.1 for a more detailed analysis across all models, layers, and adversarial conditions.
4.2 Local Results
We present the results of the local analysis applied to Mistral 7B; results for Phi3-mini-4k and LLaMA3 8B appear in Appendix C.2.
Analysis on Consecutive Layers. We computed Vietoris–Rips PH barcodes of the 2D embeddings described in Section 3.3 and extracted the total persistence and the mean birth and death of 0- and 1-bars for both clean and poisoned samples. To ensure that topological disparities are not due solely to scale differences, we repeated the analysis after normalizing activations vectors to zero mean and unit variance. Figures 10(a) and 10(b) show the average total persistence of 1-bars across 1000 activation samples. This statistic quantifies the aggregated radii of topological loops in the point clouds, capturing both their number and size. We observe clear differences between clean and poisoned activations, though these differences are diminished under normalization. Figure 10(c) shows the corresponding results for the control condition, in which activation indices are randomly permuted, disrupting meaningful neuron-wise correspondence between layers and suppressing the emergence of coherent topological features. Under this control setting, the difference in total persistence of 1-bars is significantly smaller than in both the original and normalized activation settings, suggesting that the topological signals we observe are statistically significant and not spurious artifacts.
To investigate this difference further, we measured the ratio of the total persistence of 1-bars in clean and poisoned samples (Figure 10(d)). For clean activations, this ratio steadily declines and drops below 1 near layer 12, indicating a structural phase shift. In the early layers, clean activations exhibit greater topological complexity (more or larger loops), whereas in deeper layers, poisoned activations begin to dominate in total persistence. Beyond layer 12, the trend reverses; poisoned activations accumulate longer-lived cycles, while clean activations collapse toward a simpler structure.
Poisoning appears to compress topological complexity in the early layers but, as representations propagate deeper, this constraint shifts, and the widening gap in total persistence shows that the poisoned activations increasingly diverge from the clean ones, signaling a progressively larger reconfiguration of information flow across layers. In the early layers, clean activations exhibit richer topological structure, which is gradually reduced as the network stabilizes and refines its internal representations. In contrast, poisoned activations show increasing topological complexity in deeper layers with less convergence toward stable representations. The persistent structural deviations under poisoning demonstrate that the perturbation can propagate to deeper layers, with the effect varying by whether the model refuses, executes, or ignores the injected prompt (see 13 and A.2.1). Scale-normalized activations show no such trend, confirming that these differences reflect genuine shape changes rather than mere magnitude effects.
In a real-world deployment setting, we cannot label activations as clean or poisoned in advance. To obtain a layer-specific signal without class labels, we compute the overall variance across the two classes of the total persistence of 1-bars for all activations, as shown in Figure 10(e). Empirically, this variance correlates with the absolute difference in total persistence in 1-bars between clean and poisoned activations—a quantity we used earlier as a discriminative feature. This suggests that layers with larger variance may also exhibit larger class separation. To evaluate this more rigorously, we examine whether peaks in overall variance coincide with peaks in the absolute difference in total persistence, using precision at $k$ (p@k) and permutation testing for statistical significance. The results, presented in Table 1, show a strong association between the two, with particularly more significant results for 1-bars than for 0-bars; analogous results for Mistral-7B and LLaMA3-8B appear in Table 1. Hence, high-variance layers may offer the most informative vantage points for identifying poisoned behavior when class labels are unavailable.
@@ 4.2 Local Results
Table 1: Peak analysis. Precision@$k$ for $k$=1, 3, and 5 largest peaks in total variance, and their precision in detecting the largest peaks in absolute difference between the two classes. , ∗∗ correspond to $p$-values $<$.05 and .01, respectively.
@@
Analysis on Non-Consecutive Layers. We expanded the previous analysis to activations from non-consecutive layers to show that in neighboring layers, the model operates on similar groups of neurons, leading to element-wise interactions that construct meaningful topological features distinguishing clean from poisoned datasets. However, as we examined pairs of layers that are further apart, these distinctions in interactions between clean and poisoned activations became less pronounced. Figure 11 illustrates this progression through the ratio of mean death times of 0-bars between clean and poisoned activations as the layer interval increases. For layer intervals of 1 and 3, the ratios for normalized activations and the control setting remained distinct, indicating meaningful topological interactions. However, at an interval of 10 layers, the scaled and control settings showed significant overlap, suggesting a much diminished effect of neuron interactions. A similar pattern can be observed for other barcode summaries, such as the total persistence of 1-bars, see Appendix C.2.5.
@@ 5 Discussion and Future Work
Our study provides solid evidence for PH as a useful interpretability tool for LLMs that is scalable, model-agnostic, and unifies the global organizational structure within LLM activations together with local nonlinear detail. Across six state-of-the-art models spanning diverse families and sizes, the inherent interpretability of PH barcodes reveal a consistent topological deformation under two fundamentally different attacks. This suggests a general effect in the “shape” of the representation space, which topological approaches can analyze in ways that existing methods cannot: linear probes assess linear decodability but miss latent-space structure; mechanistic interpretability traces causal pathways yet lacks a global view; and representation engineering (e.g., sparse autoencoders) uncovers local features but not topological invariants. Designing tests to compare these techniques directly with PH is therefore neither feasible nor meaningful, as they capture complementary and orthogonal aspects of model behavior. Our local, element-wise analysis of information flow within the LLM offers a novel approach to understanding the relationships between activations across layers.
Overall, our work advances both theoretical and applied perspectives in LLMs, NLP, and TDA. By demonstrating the persistent and interpretable topology and geometry of layer-wise and neuron-level interactions, it reinforces the position of topology as a powerful unifying framework for representation learning, robustness, and interpretability in neural networks papamarkou2024position .
Limitations. PH is a computationally intensive procedure: although its main kernels parallelize efficiently on modern CPUs/GPUs, the practical bottleneck is memory, since distance and boundary matrices scale roughly quadratically with the number of points. This makes exact Vietoris–Rips PH computations challenging on large high-dimensional data. To overcome this, we implemented random subsampling to compute proxy barcodes for the entire dataset, thus, our results are subject to sampling errors. However, subsampling has been well-studied in TDA; in particular, convergence results have been established chazal2014convergence ; cao2022approximating , so the sampling errors in our study are guaranteed to be bounded.
@@
Future Work. Our work inspires future investigations on whether topological compression is a general property of misalignment and how it relates to model generalization stephenson2021geometrygeneralizationmemorizationdeep ; the development of topology-aware robustness mechanisms bruelgabrielsson2020topologylayermachinelearning ; or the use of persistent Morse theory bobrowski_distance_2014 and adaptation of cycle matching approaches reani2022cycle ; fasttopologicalcyclematchinggarcia2024 to further characterize LLM representation spaces. Additionally, our analysis studied two fundamentally distinct adversarial scenarios, prompting the question of whether our topological approach is able to generalize further.
@@ Acknowledgments
[…]
@@ A.1 Persistent Homology Barcode Statistics
where $\epsilon$ is a small positive constant (e.g., $10^{-12}$) to ensure numerical stability. Higher $E$ indicates a more uniform distribution of lifetimes (no single interval dominates), whereas lower $E$ reflects a small number of long-lived intervals.
@@
We perform these computations for each barcode individually and then average over all barcodes in the same condition (locked or elicited) and (clean or poisoned).
@@
[…]
@@
Overall, poisoned conditions often yield fewer but more dominant dimension–1 features that appear later and persist longer, whereas clean spreads its total persistence across numerous shorter-lived cycles, resulting in higher entropy. Depending on the model, the specific layer where these features distinctions appear strongest differ but the nature of the difference is the same across models and layers.
@@
[…]
@@
We see similar results for Mistral 7B  with negative differences in births and deaths in earlier layers, implying that locked has larger birth/death times at those lower layers. However, the sign flips, with elicited displaying larger values for births, deaths, and persistence. Specifically, layer 32 shows a notably large positive difference (e.g., $+14.64$ for births, $+15.29$ for deaths), indicating that the final layer in elicited captures significantly later 1D cycles relative to locked. The number of 1-bars also tends to be higher in elicited at most layers, except for a minor negative at layer 23, again suggesting that elicited reveals a greater number of dimension–1 features.

### v2

@@ 4.1 Global Analysis: The Shape of Adversarial Influence
Geometric Separation of Latent States. The projection of the pruned barcode summaries over their first two principal components (Figure 7) yields a clear separation between subsamples from normal and adversarial modes across layers. This is consistent with the intuition presented in Figure 1, where a single barcode of a clean sample with $n=1000$ activations (corresponding to a point in the PCA plot) was visibly different than the barcode of a poisoned sample with same number of points. This separation signals a difference in topology between clean and poisoned subsamples; we now seek to characterize such distinction, and to test whether it is consistent across layers and models.
@@
To that end, we investigated the importance of particular features in the PCA results via a cross correlation analysis (CCA) between the pruned barcode summaries and the principal components of the PCA. CCA is a statistical method that quantifies linear relationships between two multivariate datasets by finding pairs of canonical variables with maximal correlation. The loadings are the contributions of individual features to these canonical variables, measuring their importance in capturing the relationship. We found that mean deaths of the 0-bars ranked first in all layers, and that the number of 1-bars appeared as a significant statistic as well (see Figure 15).
@@
Figure 7: PCA of pruned barcode summaries of clean vs. poisoned activations. Clear distinction appears in the two first PC projections from the PCA of the pruned barcode summaries for layers 1, 8, 16, 24, and 32. The explained variances are 0.59, 0.49, 0.52, 0.96 and 0.83, respectively.
Discriminative Power of Topological Features. We tested the power of the pruned barcode summaries in distinguishing normal and adversarial subsamples by training a logistic regression with a 70/30 split between train and test. We obtained perfect accuracy and AUC–ROC on the test data, and 5-fold cross validation over the training data (Figure 8). As a baseline comparison, we trained a linear discriminant analysis (LDA), a linear support vector machine (SVM), and a logistic regression to distinguish 1000 clean and 1000 poisoned activations, raw and after reducing dimensionality using a sparse autoencoder (AE) with hidden dimension 128; see Table 1 for results. We found that the barcode summaries outperform these methods in general, particularly for early layers. However, we emphasize that the information that they encode must be understood as complementary to that of the linear methods above, and that our true interest in the outstanding predictive power of barcode summaries resides in the fact that feature importance methods applied to the trained logistic regression allow us to interpret the differences in topology between clean and poisoned data, which is our ultimate goal.
Table 1: Comparison of predictive power with linear methods. Accuracy, with a 70/30 train/test split, of a linear discriminant analysis (LDA), a linear SVM and a logistic regression (LR) trained to distinguish 1000 raw clean activations from 1000 raw poisoned activations, with or without reducing the dimensionality of the data using a sparse autoencoder (SAE).
Layer LDA LDA (SAE) SVM SVM (SAE) LR LR (SAE)
Layer 0 0.9950 0.9875 0.8875 0.7400 0.8700 0.7400
Layer 7 1.0000 0.9975 1.0000 0.6475 0.9950 0.6350
Layer 14 1.0000 0.9975 1.0000 0.7425 0.9975 0.5875
Layer 23 1.0000 1.0000 1.0000 0.9925 1.0000 0.9575
Layer 31 1.0000 0.9975 1.0000 1.0000 1.0000 1.0000
We used Shapley (or SHAP) values to interpret the excellent performance of the regression model. Shapley values quantify the contribution (with sign) of each feature to the prediction of the model for a given input. Our analysis revealed that the mean of 0-bar deaths and the number of 1-bars strongly influence predictions, exhibiting a clear dichotomous effect: points with smaller mean death in their 0-bars and bigger number of 1-bars are typically classified as clean, whereas points with bigger mean death of their 0-bars and smaller number of 1-bars are classified as poisoned.
@@
Figure 8: Logistic regression for clean vs. poisoned activations trained on a 70/30 train/test split of the pruned barcode summaries, plotted on the projection onto the two first PCs. Accuracy and AUC–ROC on the test data and 5-fold cross validation on train data are presented for each model.
@@
The Signature of Topological Compression. Interpreting the distributions of the barcode summaries for clean vs. poisoned data reveals that adversarial conditions typically yield fewer 1-bars (loops) forming at later scales, yet persisting longer (see Figure 17). Conversely, the non-adversarial conditions tend to form earlier loops with more uniform lifetimes (higher persistent entropy). This pattern aligns with the Shapley value results (Figure 16): lower mean death times of 0-bars (i.e., more compact point clouds) are associated with predictions of “clean”, while higher values (more spread-out clouds) shift predictions toward “poisoned”. Similarly, a lower number of 1-bars tends to indicate “poisoned”, whereas a higher count suggests “clean”. In addition, a local dispersion ratio (Appendix A.2) and average cosine distance (Figure A.3) substantiated these results, revealing that adversarially influenced representation vectors become more dispersed (higher cosine distance) or concentrate variance onto specific axes (leading to flips in dispersion ratio), implying a reallocation of representational capacity toward a smaller number of large-scale features. In contrast, the non-adversarial conditions produced lower or more stable distance measures (less reconfiguration in the hidden space). Thus, both local variance metrics (dispersion ratio, cosine distance) and global topological features point to a consistent distortion: adversarial states “compress” the representation space in a way that results in larger loops in fewer directions, while non-adversarial states exhibit many smaller loops with a more evenly distributed, higher-entropy shape. This signature is robust, persisting even against adaptive attacks from the LLMail-Inject public red teaming dataset that were designed to evade activation-based defenses (see Appendix C.5). A more detailed analysis across all models, layers, and adversarial conditions is provided in Appendix A.1.
4.2 Local Analysis: Information Flow Between Layers
To investigate the fine-grained mechanisms of adversarial influence, our local analysis quantifies how information transforms between layers at the neuron level. We present the results for Mistral 7B below; see Appendix C.2 for other models.
Analysis on Consecutive Layers. Our local method revealed a structural phase shift in the network’s information flow under adversarial influence. We computed Vietoris–Rips PH barcodes of the 2D embeddings described in Section 3.3 for the raw activations; their normalization to zero mean and unit variance, to ensure that topological signals are not due solely to scale differences; and a control condition where neuron indices are randomly permuted, disrupting any neuron-wise correspondence between layers. We measured the topological complexity by the total persistence of 1-bars, and found significant differences across layers in the raw and normalized activations (Figures 9(a) and 9(b)). In addition, as shown by the ratio of topological complexity between clean and poisoned activations (Figure 9(d)), clean inputs initially exhibit a more complex structure that simplifies in deeper layers. In contrast, poisoned activations start simpler but their topological complexity increases, diverging significantly from the clean activations around layer 12. This suggests that adversarial influence causes a major reconfiguration of information processing in the model’s deeper layers. The signal’s disappearance in the permuted control condition (Figure 9(c)) confirms that this effect is tied to specific neuron-to-neuron pathways and is not a statistical artifact.
In a real-world setting without labels, these informative layers can still be identified. We found that the overall variance of a topological feature across all samples strongly correlates with the magnitude of the clean-vs.-poisoned difference (Figure 9e). As shown in Table 2, via precision at $k$ (p@k) and permutation testing for statistical significance, layers with the highest variance are reliable indicators of layers with the largest class separation, providing a practical, unsupervised signal for locating where adversarial effects are most prominent.
@@ 4.2 Local Analysis: Information Flow Between Layers
Table 2: Peak analysis. Precision@$k$ for kk=1, 3, and 5 largest peaks in total variance, and their precision in detecting the largest peaks in absolute difference between the two classes. , ∗∗ correspond to $p$-values <<.05 and .01, respectively.
@@
Analysis on Non-Consecutive Layers. We expanded the previous analysis to activations from non-consecutive layers to show that in neighboring layers, the model operates on similar groups of neurons, leading to element-wise interactions that construct meaningful topological features distinguishing clean from poisoned datasets. Figure 10 illustrates the ratio of the mean death times of 0-bars between clean and poisoned activations as the layer interval increases. For layer intervals of 1 and 3, the ratios for normalized activations and the control setting remained distinct, indicating meaningful topological interactions. However, at an interval of 10 layers, the scaled and control settings showed significant overlap, suggesting a much diminished difference in the interactions in clean and poisoned data. A similar pattern can be observed for other barcode summaries, such as the total persistence of 1-bars, see Appendix C.2.5.
@@ 5 Discussion and Future Work
Our global and local analyses provide converging evidence for a fundamental principle, where adversarial influence manifests as “topological compression” of an LLM’s latent space. This behavior—a shift from compact, diverse structures to more dispersed, topologically simpler ones—is a consistent, architecture-agnostic phenomenon that holds across different model architectures, sizes, and attack vectors. This topological approach offers a distinct and complementary form of interpretability that is relational rather than compositional. While methods such as sparse autoencoders (SAEs) (Cunningham et al., 2023) are powerful for identifying the “building block” features of a representation, they analyze each activation in isolation. This makes them inherently blind to the nonlinear, relational geometry that emerges from the interactions between activations. Furthermore, because the feature dictionaries learned by SAEs are specific to a single set of model weights, they cannot be reliably compared across different models or fine-tuning stages. Our PH-based framework circumvents these limitations by computing intrinsic, coordinate-free geometric properties, providing a stable basis for comparison and enabling a comprehensive characterization of the shape of adversarial influence.
The implications of our work extend to the core of interpretability and AI safety. Our findings contribute to a growing body of evidence that a model’s behaviors are encoded in the geometry of its latent space. This perspective aligns with work showing that memorization corresponds to a reduction in the effective dimensionality of the representation manifold (Stephenson et al., 2021), and that the success of linear probes may stem from their ability to approximate more complex topological structures (Engels et al., 2025b). Our discovery that adversarial influence induces a “topological compression” provides new evidence for this hypothesis, suggesting that a collapse in geometric complexity is a quantifiable signature of out-of-distribution states. Our findings reframe key safety properties such as robustness not merely as abstract behavioral outcomes, but as measurable characteristics of the representation space itself.
Limitations. The primary limitation of our study is the memory requirements of PH, as the distance and boundary matrices required for exact Vietoris–Rips computations scale quadratically with the number of points. To manage this on our large datasets, we implemented random subsampling, which is well-studied in TDA with established convergence results ensuring that the sampling errors in our study are bounded (Chazal et al., 2014; Cao and Monod, 2022).
@@
Future Work. Our study opens several avenues for future investigation, such as exploring whether topological compression is a general property of model misalignment (Stephenson et al., 2021); developing topology-aware robustness mechanisms (Brüel-Gabrielsson et al., 2020); applying persistent Morse theory (Bobrowski and Adler, 2014); and adapting cycle matching approaches (Reani and Bobrowski, 2022; García-Redondo et al., 2024) to further characterize LLM representation spaces. Further study is also needed to see if these topological signatures generalize to an even broader range of adversarial scenarios.
@@ Acknowledgments
A.F. wishes to thank Daniel Jones for helpful discussions. I.G.R. is funded by a London School of Geometry and Number Theory–Imperial College London PhD studentship, which is supported by the EPSRC grant No. EP/S021590/1. Q.W. is funded by a CRUK–Imperial College London Convergence Science PhD studentship, which is supported by Cancer Research UK under grant reference CANTAC721\10021 (PIs Monod/Williams). H.D. is supported by the research program “Change is Key!” supported by Riksbankens Jubileumsfond (under reference number M21-0021). H.D. and A.M. are supported by the EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
@@ A.1 Persistent Homology Barcode Statistics
where $\epsilon$ is a small positive constant (e.g., $10^{-12}$) to ensure numerical stability. Higher $E$ indicates a more uniform distribution of lifetimes (no single interval dominates), whereas lower $E$ reflects a small number of long‐lived intervals.
@@
We perform these computations for each barcode individually and then average over all barcodes in the same condition (elicited or elicited) and (clean or poisoned).
@@
Table 3: Dimension-1 persistent homology differences ($\text{clean}-\text{poisoned}$) in key metrics for three models across several layers. Positive values mean the clean condition has a higher value, while negative indicates poisoned is higher for that metric. All entries rounded to four decimals.
@@
Overall, poisoned conditions often yield fewer but more dominant dimension-1 features that appear later and persist longer, whereas clean spreads its total persistence across numerous shorter-lived cycles, resulting in higher entropy. Depending on the model, the specific layer where these features distinctions appear strongest differ but the nature of the difference is the same across models and layers.
@@
Table 4: Dimension-1 persistent homology differences ($\text{elicited}-\text{locked}$) for two models across multiple layers. Positive values indicate that the elicited condition has higher values; negative means locked is higher for that metric.
@@
We see similar results for Mistral 7B  with negative differences in births and deaths in earlier layers, implying that locked has larger birth/death times at those lower layers. However, the sign flips, with elicited displaying larger values for births, deaths, and persistence. Specifically, layer 32 shows a notably large positive difference (e.g., +14.64+14.64 for births, +15.29+15.29 for deaths), indicating that the final layer in elicited captures significantly later 1D cycles relative to locked. The number of 1-bars also tends to be higher in elicited at most layers, except for a minor negative at layer 23, again suggesting that elicited reveals a greater number of dimension-1 features.